### v1

# Dissipative spatiotemporal soliton in a driven waveguide laser

Abstract
A distributed Kerr-lens mode locking regime can be realized in a waveguide laser by spatial profiling of the pump beam, thus creating a spatio-temporal soliton. Additional slow temporal modulation of the pump source stabilizes the spatio-temporal solution in a broad range of parameters, which are defined by the dynamic gain saturation. We choose a Cr:ZnS waveguide laser as a practical example, but such a regime is feasible in various waveguide and fiber oscillators. A far-reaching analogy with Bose-Einstein condensates allows using this approach to stabilization of the weakly dissipative BECs.
journal: jocn
1 Introduction
The unprecedented progress in the field of mesoscopic physics in its general context, which comprises control of macro-quantum states like Bose-Einstein condensate (BEC) [1, 2, 3, 4] and quantum supersolids [5], plasma and turbulence [6, 7, 8], attosecond [9], relativistic and high-energy physics [10, 11, 12, 13], metaphorical modeling and quantum computations [14, 15, 16, 17], physics of ultra-strongly coupled light-matter systems [18], nano-processing [19, 20, 21], and other ranges of science, medicine, and technology, became feasible due to the growth understanding of strongly nonlinear phenomena in far-from-equilibrium systems [22, 23]. The coherent, self-organized, and well-localized patterns, namely solitons, play a crucial role in most of these phenomena. The class of such structures, dissipative solitons (DSs), is particularly important because DS develops in open systems, where energy exchange with an environment defines its integrity, stability, and coherence [24, 25]. In photonics, DSs are a road to generating ultrafast, robust, and energy-scalable pulses in mode-locked lasers [26, 27] bringing a high-field physics on tabletops of a mid-level university lab [12, 28].
The modern tendencies in generating and exploring laser DSs are based on a controllable enhancement of self-organizing effects induced by nonlinearities and affected both temporal and spatial degrees of freedom [23]. A partial realization of this concept is a distributed Kerr-lens mode-locking (DKLM) [29, 30, 31]. In a solid-state laser, this technique uses a nonlinear medium that affects a laser beam via self-focusing, thereby changing an effective gain in a laser resonator [32]. The transversal spatial structure of a field in this type of device is close to a fundamental mode of laser oscillator and, thus, trivial, so only longitudinal modes need synchronization. A remarkable breakthrough in energy scalability of such oscillators was demonstrated in both anomalous and normal group-delay dispersion regimes (ADR and NDR, respectively) [33, 34, 30]. It was found that such broad-range scalability could be explained by energy out/in-flows induced by the nontrivial phase structure of DS [27, 35], and the underlying theory of the DS resonance (DSR) [36] were developed.
Using DS in the NDR fiber lasers [37] exploits the enhanced and well-controllable nonlinearities. That bridges fiber and solid-state ultrafast laser photonics [23]. The keystone is utilizing a nonlinear propagation of many interacting spatial modes (e.g., in the so-called multimode fibers, MMF) [38, 39]. Such a nonlinear spatial mode coupling could be caused by nonlinear refraction (or attractive boson interaction in BEC [4]) characterized by the coefficient $n_{2}$. A “confining potential” defined by spatial-dependent refractive index $n(x,y)$ (“graded refractive index fibers,” or GRIN, nonlinear lattices [40, 41], or laser-induced confinement in BEC) enhances the relaxation of the higher-order modes to a ground-state, i.e., lowest-order spatial mode.
In this work, we propose utilizing a nonlinear mode condensation for DKML aimed at a spatio-temporal mode-locking (STML) [42, 43].
Such an approach utilizes a dissipative (i.e., $\Im[n(x,y)]\neq 0$, where $n$ is, e.g., a complex index of refraction in optics) trapping potential [44] for a spatiotemporal DS (STDS) without involving other nonlinear dissipative processes. The 2D dissipative potential, unlike that in Refs. [45, 46] transversal $(x,y)$-confinement induced by a complex $n(x,y)$ can be associated with a so-called “cigar-type” trapping potential in a weakly dissipative BEC or GRIN (see Table 1) [15]. Then, an evolutionary (“slow”) coordinate $T$ is associated with time for BEC or propagation distance $Z$ in a fiber or cavity round-trip in a laser.
A further step is to introduce “longitudinal” (“slow” time $T$-dependent, or propagation length $Z$-dependent) modulation of $\Re[n(x,y,T)]$ (or $\Re[n(x,y,Z])$, which is one of the management techniques stabilizing an ST soliton in BEC [47]. In a mode-locked laser, it is also possible to identify a “fast” coordinate (local time $t$) associated with the frame co-moving with a soliton. The dependence of trapping potential on that coordinate allows a 3D confinement corresponding to a “pancake-like” potential in BEC. In a laser, this corresponds to an active mode-locking driven by an external synchronous phase modulation[48, 49, 50], which varies as $\Re[\Delta n(t)]\propto\cos(2\pi t/T_{R})$, where $T_{R}$ is a cavity round-trip period. This modulation is slow (pulse duration $\tau_{p}$ is much shorter than $T_{R}$) and can be described by $\Re[\Delta n(t)]\propto t^{2}+\mathrm{H.O.T.}$. It was found that such a modulation stabilizes a driven cavity DS [51, 52] and STDS [53]. Here, we intend to exploit a concept of 3D dissipative confinement to develop a spatiotemporal mode-locking mechanism to generate stable and energy-scalable STDS (“light bullet”). This technique can be considered as a road to forming a stable mass-scalable STDS in BEC and pattern formation in dissipative turbulent systems, which could be modeled by the dissipative version of the Gross-Pitaevskii equation (GPE) [4, 54]. In contrast to [55, 53], a transversely graded dissipation (i.e., $\Im[n(x,y)]\neq 0$) will be supplemented by dissipative trapping along $t$-coordinate ($\Im[n(t)\neq 0$), which is a result of periodical loss modulation (or synchronous pumping) synchronized with the laser cavity round-trip. For BEC, this corresponds to a “pancake-like” dissipative confining potential.
Table 1: Correspondences between photonics and BEC [15]
Laser
BEC
Propagation coordinate $Z$
$\leftrightarrow$
time $T$
Pulse local time $t$
$\leftrightarrow$
third spatial coordinate $z$
Diffraction + anomalous GDD
$\leftrightarrow$
boson kinetic energy
Kerr-nonlinearity
$\leftrightarrow$
boson attractive colliding potential
GRIN
$\leftrightarrow$
cigar-like ($x$, $y$) trapping potential
Non-selective loss
$\leftrightarrow$
homogeneous BEC dissipation
Graded gain + local time modulation
$\leftrightarrow$
pancake ($x,y,z$)-shaped inflow of bosons
Spectral dissipation
$\Leftrightarrow$
“kinetic cooling”
2 A “soft-aperture” DKLM laser
2.1 Model
Let us consider a waveguide laser composed of a fiber or crystalline waveguide active medium enclosed in a Fabri-Perót or ring resonator and driven by a pump beam with a parabolic-like transverse $G(x,y)-$ profile and a pump intensity varying with time $t$ at the frequency synchronized with a laser repetition rate. The $t-$coordinate corresponds to a “local time” (in a coordinate frame co-moving with an STDS) in photonics or to a longitudinal space coordinate for BEC (see Fig. 1 and Table 1) [15].
Figure 1: Multi-scale scheme with a local-time coordinate $t$ synchronized with an STDS group velocity and a global evaluation coordinate $Z$ (see Table 1). The gain (filled by a red) is profiled along the radial coordinate $r$ and the local time $t$. The last is synchronized with a laser round-trip period.
Such pump beam acts as a “soft aperture” guiding the laser beam and controlling (“cleaning”) its mode structure so that a laser field $a(z,x,y,t)$ tends to condensate into a lowest-order mode. Simultaneously, $t-$modulation provides confinement in the time domain. Thus, this scheme corresponds to a three-dimensional (“pancake-like”) dissipative confinement for a BEC or a hybrid mode-locking under synchronous pumping or an active assisting mode locking by synchronous loss modulation [56, 57, 50, 52].
For modeling a (3D+1)-dimensional trapped STDS, we use a distributed model describing the evolution of the $(z,x,y,t)$-dependent field $a$ propagating along the $z$-axis (see Table 1 for comparison with BEC) of an active waveguide with the linear and nonlinear refractive indexes $n(x,y)$ and $n_{2}$, respectively. Under the action of $(x,y,t)$-graded pump (Fig. 1), the evolution can be described by a sort of the driven nonlinear Helmholtz equation:
$-\left[\frac{1}{2k_{0}n_{0}}\left(\Delta_{x,y}-\frac{\partial^{2}}{\partial z^{2}}\right)+(\frac{\beta_{2}}{2}-i\tau)\frac{\partial^{2}}{\partial t^{2}}\right]a=$ (1)
$\left[\frac{k_{0}}{2n_{0}}\left(n(x,y)^{2}-n_{0}^{2})\right)+i\left(L-G(x,y,t)\right)+k_{0}n_{2}\left|a\right|^{2}+i\frac{\partial}{\partial z}\right]a,$
where $\Delta_{x,y}$ is a Laplace operator in the Cartesian coordinates, $n_{0}$ is a waveguide cladding refractive index, $k_{0}$ is a free-space wave number, and the second-order derivative over $z$ contributes only for large beam numerical apertures NA [58, 59]. The terms in Eq. 1 define: (I) – diffraction, (II) – group-delay dispersion with the coefficient $\beta_{2}$, (III) – spectral dissipation with the squared inverse bandwidth $\tau$, (IV) – trapping potential induced by GRIN, (V) – $(x,y,t)$- graded saturated gain with the coefficient $G$ and linear net-loss $L$, and (VI) – Kerr-nonlinearity. One has to emphasize the decisive contribution of a saturable and transversely profiled gain $G(x,y,t)$ in Eq. (1).
For a low NA, one may omit the $\frac{\partial^{2}}{\partial z^{2}}$-term in Eq. (1). That leads to the nonlinear driven Schrödinger, or the Lugiato-Lefever equation (LLE), which is a test bed model for the study of solitons and non-equilibrium patterns in driven Kerr resonators, micro-cavity and fiber lasers, VCSEL, etc. [60, 61, 62, 63].
The rescaling $\psi=a\exp(-iV_{0}z)$ ($V_{0}=(k_{0}/2n_{0})(n_{1}^{2}-n_{0}^{2})$, and $n_{1}$ is a waveguide core refractive index) leads to the dimensionless mean-field LLE [64, 65, 66]:
$i\frac{\partial\psi\left(Z,X,Y,t^{\prime}\right)}{\partial Z}=\left\{-\frac{1}{2}\left[\frac{\partial^{2}}{\partial X^{2}}+\frac{\partial^{2}}{\partial Y^{2}}+\frac{\partial^{2}}{\partial t^{\prime 2}}-2i\tau\frac{\partial^{2}}{\partial t^{\prime 2}}\right]\right\}\psi+$
$\left\{(X^{2}+Y^{2})-\left|\psi\right|^{2}-i\left[\Lambda+\kappa(X^{2}+Y^{2})+\omega t^{\prime 2}\right]\right\}\psi,$ (2)
for the normalizations shown in Table 2 [64, 67].
The dimensional extension by the “fast-time” (or “local time”) coordinate $t^{\prime}$ allows describing the pulse dynamics in the co-moving frame $(Z,X,Y,t^{\prime})$. Since an ultrashort pulse is sensitive to a group-delay dispersion (GDD), it has to be taken into account by including the $\beta_{2}$-term in Eq. (1). The normalized version (2) assumes an anomalous net-GDD.
In this work, we assume the parabolic-like transverse profile of the gain beam with the size. Such transversely confined nonlinear gain acts as a “soft aperture” with an effective size $\approx w_{p}\sqrt{G_{0}-\Lambda}$ (Table 2). Our model introduces a gain modulation along the $t-$axis described in a parabolic approximation by the modulation coefficient $\omega$ [50, 57]. The last resembles the mechanism of the so-called active mode-locking (AM) [68]. In contrast to [53], the proposed method does not use additional phase modulation but only a pumping pulse train synchronized with a resonator period. For simplicity, we manipulate with a graded saturated gain $G(X,Y,t)=G(0,0,0)\times\left[1-\kappa(X^{2}+Y^{2})-\omega t^{\prime 2}\right]$, where $G(0,0,0)$ or $G_{0}$ (see Table 2) corresponds to a saturated gain along a waveguide axis at $t^{\prime}=0$, and $\kappa=G_{0}w_{p}^{-2}$, $\omega=G_{0}\mathscr{T}^{-2}$ ($\mathscr{T}$ is a “pump pulse duration”). Such an approximation is valid under the condition of quasi-steady state operation of a laser in the vicinity of the lasing threshold, where the saturated net-loss coefficient $\Lambda^{\prime}=L-G_{0}$ is close to zero [55].
Another essential factor in Eqs. (1,2) is a spectral dissipation described by the parameter $\tau$. This parameter is inversely proportional to a squared spectral width of the gain band or spectral filter (Table 2).
Thus, one may treat the resulting equation (2) as a dissipative extension of the driven Gross-Pitaevskii equation, which is a well-known tool for analyzing trapped BEC (Table 1) [4].
Table 2: Normalizations for the LLE (2)
Value
Normalizations and definitions
$(X,Y)$
transverse coordinates, $(x,y)/w_{T}$
$Z$
propagation distance, $z/\zeta$
$t^{\prime}$
local time, $t/\sqrt{|\beta_{2}|\zeta}$
$\zeta$
propagation scale, $k_{0}n_{o}w_{T}^{2}$
$w_{T}$
transverse scale, $4]{w_{0}^{2}/k_{0}^{2}n_{0}\delta}$
$\delta$
GRIN contrast, $n_{1}-n_{0}$
$\left(k_{0}/2n_{0}\right)\left(n_{1}^{2}-n(x,y)^{2})\right)$
GRIN confining potential,  $\simeq k_{0}\left(n_{1}-n(r)\right)$
$\left|\psi\right|^{2}$
intensity, $k_{0}n_{2}\zeta\left|a\right|^{2}$
$\Lambda^{\prime}$
net-loss, $\Lambda-G_{0}$
$\Lambda$
averaged loss, $LL_{w}/\zeta$
$\kappa$
spatial confinement parameter, $G_{0}/w_{p}^{2}$
$\omega$
temporal confinement parameter, $G_{0}/\mathscr{T}^{2}$
$w_{p}\sqrt{G_{0}-\Lambda}$
soft aperture size
$\tau$
squared inverse spectral filter, bandwidth normalized to $|\beta_{2}|\zeta$
$k_{0}$
wave-number, $2\pi/\lambda$
$\lambda$
central wavelength
$n_{1}$
GRIN core refractive index
$n_{0}$
GRIN cladding refractive index
$n_{2}$
nonlinear refractive index
$L$
loss coefficient
$L_{w}$
waveguide (“lattice”) length
$G_{0}$ or $G(t=0,x=0,y=0)$
maximum saturated gain coefficient
$w_{p}$
pump beam size
$\mathscr{T}$
pump pulse width
2.2 Variational approximation
The study of STDS of Eq. (2) is based on analytical and numerical approaches. The first uses the well-known variational approximation (VA) [55]. The generating Lagrangian for a non-dissipative part of (2) is
$\mathscr{L}=\frac{i}{2}\left(\psi^{*}\partial_{Z}\psi-\psi\partial_{Z}\psi^{*}\right)+\frac{1}{2}\left(\left|\partial_{X}\psi\right|^{2}+\left|\partial_{Y}\psi\right|^{2}+\left|\partial_{t^{\prime}}\psi\right|^{2}\right)-$
$\left|\psi\right|^{4}+\left(X^{2}+Y^{2}\right)\left|\psi\right|^{2}.$ (3)
The dissipative factors can be described in the form of a “force” source term
$\mathcal{Q}=-i\left[\Lambda^{\prime}+\kappa(X^{2}+Y^{2})+\omega t^{\prime 2}-\tau\frac{\partial^{2}}{\partial t{{}^{\prime 2}}}\right]\psi$ (4)
in the Euler-Lagrange equation
[…]
in agreement with Kantorovitvh’s method (see [55, 15, 53] and references herein).
It is convenient to assume the axial symmetry, which corresponds to the symmetry of a cylindrical waveguide, and transit to a radial coordinate $R=\sqrt{X^{2}+Y^{2}}$. The next assumption is a Gaussian-sech approximation for an STDS profile:
$\psi(Z,R,t^{\prime})=\alpha(Z)\operatorname{sech}\left[\frac{t^{\prime}}{\Upsilon(Z)}\right]\times$
$\exp\left[i(\phi(Z)+\chi(Z)t^{{}^{\prime}2}+\theta(Z)R^{2})-\frac{R^{2}}{2\rho(Z)^{2}}\right],$ (6)
where the $Z-$dependent STDS parameters of f in Eq. (5) are: $\alpha$ amplitude, $\phi$ phase, $\chi$ and $\theta$ are the temporal and spatial chirps, respectively; $\rho$ and $\Upsilon$ are the beam size and STDS duration, respectively. After some algebraic manipulations, one can obtain the expressions for the steady-state (i.e., $Z-$independent) STDS parameters (the phase $\phi$ is irrelevant in our context):
$\rho^{2}=\frac{1}{384\pi^{2}\kappa\tau\Upsilon^{2}}\times$ (7)
$\biggl{[}15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}-15\right)-$
$64\tau\left(15\tau+2\pi^{2}\left(\tau+3\Lambda^{\prime}\Upsilon^{2}\right)\right)\biggr{]},$
$\chi=\frac{\sqrt{\Upsilon^{4}\left(64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225\right)}-15\Upsilon^{4}}{16\pi^{2}\tau\Upsilon^{4}},$ (8)
$\alpha_{0}^{2}=\frac{-6\kappa^{2}\rho^{8}-6\rho^{4}+6}{2\rho^{2}},$
$\theta=-\frac{\kappa\rho^{2}}{2}.$
The remaining equation for the STDS duration $\Upsilon$
$144\left(\pi^{2}-25\right)\tau^{2}+\left(16\left(3+\pi^{2}\right)\tau^{2}+45\right)\times$ (9)
$\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}+96\pi^{4}\Upsilon^{4}\omega+$
$+\frac{2^{13}3^{3}\pi^{4}\kappa\tau^{3}\Upsilon^{4}\left(-\frac{A^{4}}{3\times 2^{15}\pi^{8}\kappa^{2}\tau^{4}\Upsilon^{8}}-\frac{3A^{2}}{2\pi^{4}\kappa^{2}\tau^{2}\Upsilon^{4}}+1\right)}{A}=675,$ (10)
$A=64\tau\left(2\pi^{2}\left(3\Lambda^{\prime}\Upsilon^{2}+\tau\right)+15\tau\right)-$
$15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}-15\right)$
must be solved numerically.
2.3 STDS parameters
Figs. 2-5 show the STDS width $\Upsilon$ and peak intensity $\alpha_{0}^{2}$ obtained from Eqs. (7-9). $\Upsilon$ grows, and $\alpha_{0}^{2}$ decreases with $\kappa$ when a spectral dissipation $\tau$ is sufficiently low (Figs. 2, 3). Such a tendency corresponds to that of a DKLM regime with $\omega=0$ and manifests the growth of spatial mode loss with an aperture squeezing [55].
In the vicinity of a “threshold” ($\Lambda^{\prime}\to$0), the additional longitudinal (3D) trapping defined by $\omega$ noticeably squeezes an STDS with the simultaneous peak power growth (Fig. 3). The additional solution branch in the region of small $\kappa$ (Fig. 2) disappears with $\Lambda^{\prime}\to$0. In general, the STDS intensity growth and its width shortening with $\omega$ could result from an appearance of the external confinement along the $t-$ coordinate.
The spectral dissipation increases with $\tau$. That leads to an STDS collapse for $\kappa\to|\Lambda^{\prime}|$ when the net-loss coefficient $\Lambda^{\prime}$ is more significant in comparison with Fig. 3. Such a collapse is asymmetrical in space and time. That is, STDS spreads along $t$ and squeezes on $R$. A longitudinal mode selection growth with $\tau$ (i.e., STDS spreading in the time domain) parallels a transverse mode selection (STDS spatial squeezing) in the spatial domain.
The solutions bifurcate close to the threshold $\Lambda^{\prime}$ (Fig. 5). Such behavior leads to the thought of stability loss here. These oddities require numerical simulations of STDS dynamics presented below.
Figure 2: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=$-0.01, $\tau=$0.01.
Figure 3: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.001, $\tau=$0.01.
Figure 4: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.01, $\tau=$0.1.
Figure 5: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ (b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.001, $\tau=$0.1.
2.4 STDS stability
We performed large-scale numerical simulations of Eq. (2) based on the COMSOL Multiphysics software to analyze the stability of the analytical solutions based on the VA. The characteristic case is illustrated in Fig. 6, where the numerical STDS profile corresponds to one of the VA solutions from Fig. 4. The red line shows the analytical and numerical solutions shown by the black curve. One can see an oscillatory approach from the last to the first, analogous to the damping of excitations in a weakly dissipative BEC [69]. Interestingly, the peak intensity of the numerical solution is slightly lower than the analytical one. The difference in the exact solution from the Gaussian ansatz for VA could explain it.
[…]
Besides the stable STDS solutions, the numerical simulations demonstrate three main destabilization scenarios. The first is the STDS decay with the $|\Lambda^{\prime}|$-decrease. The second is the collapse-like behavior for small $\kappa$, $\omega$, and $\tau$. Both demonstrate the importance of dissipative factors (i.e., dissipative confinement and spectral filtering) for STDS stability.
The third scenario is the DS splitting with possible oscillations between multiple pulse states. Fig. 7 shows such regimes. The splitting in a local time domain could be explained by weaker confinement in the time domain due to lower $\omega$. However, a more precise analysis of the STDS requires considering the gain saturation dynamics.
Figure 7: Contour plot of the STDS peak intensity $\alpha_{0}^{2}$ at $Z=$40000 (a) and 35000 (b) for $\omega=10^{-5}$ (a) and $2\times 10^{-6}$ (b), $\kappa=$0.005, $\Lambda^{\prime}=$-0.01, and $\tau=$0.1.
2.5 STDS stability under the saturable gain dynamics
To take into account the dynamical gain saturation, it is convenient to use the Frantz-Nodvik equation [70]:
$\frac{\partial G_{0}}{\partial t}=\frac{G_{ss}-G_{0}}{T_{r}}-\frac{G_{0}\left|\psi\right|^{2}}{E_{s}},$ (11)
where $G_{ss}$ is a small signal gain, $T_{r}$ is a gain relaxation time, $E_{s}=h\nu/S\sigma_{em}$ is a gain saturation energy ($\nu$ is a central gain-band wavelength coinciding with the STDS carrier frequency, $S$ is a beam area, $\sigma_{em}$ is a stimulated emission cross-section). Specifically, we base on the characteristics of the Cr:ZnS waveguide laser [71]. Table 3 presents the corresponding dimensional laser parameters.
Table 3: Correspondences between photonics and BEC
Cr:ZnS laser parameters (see Table 1) values
gain relaxation time $T_{r}$, $\mu s$ 4.9
[…]
central wavelength $\lambda$, $\mu$m 2.3
$w_{0}$, $\mu$m 40
$n_{0}$ 2.2629
$\delta$ 0.0077
numerical aperture $NA=\sqrt{n_{1}^{2}-n_{0}^{2}}$ $\approx 0.19$
$w_{T}$, cm 0.001
$\zeta$, cm 0.03
[…]
The calculations demonstrate a visible effect of the gain saturation on the DS dynamics. For a small level of linear loss $\Lambda$, some characteristic profiles of STDS are shown in Fig. 8. As one can see, a sufficiently large $\omega$ stabilizes DS ($a$), but the driving frequency decrease leads to the STDS shape distortion, which is periodically asymmetrical along $t$-coordinate. Oscillations accompany these distortions and result in STDS decay. STDS cannot be stabilized for the parameters considered for a smaller $\Lambda$ (Fig. 5).
[…]
The region of STDS stability is shown in Fig. 9 for a larger linear loss coefficient $\Lambda$. The STDS decay confines this region for larger $\kappa$, so there are no stable pulses for $\kappa\gtrsim 0.006$. The difference with the analytical results lies in the reduction of a net gain required for the soliton formation. Such a gain reduction is a direct consequence of the gain saturation. For smaller $\kappa$, the main destabilization factor is periodical pulse splitting, which alternates with a single-pulse regime (Figs. 8 (b-d) and inset in Fig. 9). The asymmetry of the STDS complex appears periodically due to dynamic gain saturation. When $\kappa$ approaches zero, very strong STDS power and width oscillations appear. They reach more than two orders of magnitude, so STDS is periodically “reborn”. This phenomenon confines the stability region from below on $\kappa$.
Figure 9: Region of STDS stability on ($\kappa$$\omega$)-plane under the action of dynamical gain saturation. $\Lambda=$0.1, and other parameters as in Fig. 8. Inset shows the $|\alpha_{0}|^{2}$ contour plot for $\kappa=5\times 10^{-4}$, $\omega=\times 10^{-6}$.
3 Discussion and Conclusion
Similar to the mechanism of STDS stabilization by non-dissipative confinement along the longitudinal ($t$) coordinate [53], which corresponds to an additive phase mode-locking in a laser, the dissipative confinement due to synchronous pulse pumping or active amplitude mode-locking leads to a stable STDS, as well. Such stabilization is possible at very low values of the modulation “depth” parameter $\omega=G_{0}/\mathscr{T}^{2}\approx\Lambda/\mathscr{T}^{2}\approx 10^{-4}$. This means that gain variations along a laser axis could be very small. For instance, it means that a pumping pulse, synchronized with the cavity round-trip, could be very “long” compared to the STDS temporal width. The contribution of dynamical gain saturation lowers this value to $\approx 10^{-6}$. Simultaneously, there is a lower threshold on the value of the saturated net gain $|\Lambda^{\prime}|$. In our case, it was $\approx 0.01$. Too low net gain could not provide the 3D dissipative confinement and, thereby, STDS stabilization. We also found that the most optimal stabilization region is located near $\kappa\approx|\Lambda^{\prime}|/2$. Considering the dynamical gain saturation, the corresponding value is $\kappa\approx\Lambda/4-\Lambda/3$. It is important that spectral dissipation ($\tau$) is required for STDS stabilization. One may assume that the too-low value of this parameter cannot prevent a tendency to the STDS collapse.
The numerical simulations demonstrated that, besides the STDS destabilization due to decay for low $|\Lambda^{\prime}|$, $\kappa\rightarrow 0\land|\Lambda^{\prime}|$, and small $\omega$ and $\tau$, there is a tendency to STDS splitting in the time domain with the growth of $\kappa$ if $\omega$ is not sufficiently high. That means a “transverse” confinement needs an adequate “longitudinal” one. Despite the case of longitudinal phase confinement [53], we did not observe complicated spatial dynamics and splitting that testifies to the intensification of the mode cleaning process. Although “ring-like” spatial structures and the STDS “breathing” in the transverse direction are possible in the initial stages of evolution.
We found that the gain saturation could significantly impact the STDS dynamics. For the considered values of gain coefficient $G_{ss}$ there is a confined range of driving frequency $\omega$ providing the STDS stability, and this region is the broadest in the vicinity of $\kappa\approx 0.0003-0.0004$. The main mechanism of destabilization is STDS decay. A possible explanation is the following. Too small $\omega$ does not provide sufficient confinement along $t-$dimension for a chosen $G_{ss}$. Too large driven frequency squeezes a gain window so STDS cannot be amplified sufficiently to start a Kerr effect for spatial squeezing, and spatial confinement becomes destructive. Also, the dynamical gain saturation defines some maximum $\kappa$ because it decreases an effective gain, and the spatial losses lead to the STDS decay. STDS peak power and width oscillations appear for small $\kappa$. These oscillations can be extremely strong and accompanied by alternating between asymmetric one and two pulses.
We must note, that the cavity round-trip period could excess essentially $2L_{w}/c$ in order to provide the STDS energy scalability [59]. This method is practiced in high-energy passively mode-locked oscillators (e.g., see [12, 30]). In a waveguide oscillator, it is possible to use the Herriott cell or fiber loop for this aim. The last approach could be the most effective for the STDS generation in a fiber laser [53].
In conclusion, we propose a mechanism of STDS stabilization using longitudinally and transversely graded gain in a waveguide laser with transversely graded refractive index and Kerr nonlinearity. Practically, that means realizing a DKLM in a waveguide laser operating in ADR and pumped by a Gaussian beam with slow amplitude temporal modulation synchronized with a laser repetition rate. In some sense, that is a dissipative analog of “pancake-like” confining potential in a weakly dissipative BEC. The analytical solutions for such STDS are obtained in the framework of VA depending on the parameters of the pump beam, such as its width, modulation depth, and saturated net gain. The essential ingredient of our model is spectral dissipation due to finite spectral gain bandwidth, which contributes to STDS stabilization. Numerical simulations under cylindrical symmetry conditions demonstrate the STDS’s stability within a limited region of parameters defined by net gain, pump beam width, and spectral bandwidth. Considering the dynamical gain saturation, we defined the main destabilization scenarios: decay and STDS oscillations asymmetrically alternating one and two pulses. We assume that the self-control of the transverse modes by the combination of enhanced nonlinearity, specific for the distributed nonlinear systems like waveguides composed from highly-nonlinear media, with the dissipative confinement, could provide the breakthrough in developing the new generation of high-stable and energy scalable sources of STDS. In principle, this approach is a further development of the DKLM concept that proved its efficiency for energy-scalable solid-state lasers. Moreover, it could be useful for stabilizing other localized coherent structures, e.g., a weakly dissipative BEC.
Funding
The Norwegian Research Council projects #303347 (UNLOCK), #326503 (MIR), #32641 (Lammo-3D), and ALTA Lasers AS supported this work.

### v2

@@
 A distributed Kerr-lens mode locking regime can be realized in a waveguide laser by spatial profiling of the pump beam, thus creating a spatio-temporal soliton.
-Additional slow temporal modulation of the pump source stabilizes the spatio-temporal solution in a broad range of parameters, which are defined by the dynamic gain saturation.
+Additional slow temporal modulation of the pump source stabilizes the spatio-temporal solution in a broad range of parameters defined by the dynamic gain saturation.
 We choose a Cr:ZnS waveguide laser as a practical example, but such a regime is feasible in various waveguide and fiber oscillators.
-A far-reaching analogy with Bose-Einstein condensates allows using this approach to stabilization of the weakly dissipative BECs.
-journal: jocn
+A far-reaching analogy with Bose-Einstein condensates allows this approach to stabilize the weakly dissipative BECs.
+keywords:
+waveguide laser , spatiotemporal soliton , mode-locking , mid-infrared lasers
+journal: Optics Communications
+[…]
+[first]organization=Department of Physics, Norwegian University of Science and Technology, addressline=, city=Trondheim, postcode=7491, country=Norway
+[…]
+[second]organization=Institut für Photonik, TU Wien, addressline=Gusshausstrasse 27/387, city=Vienna, postcode=1040, country=Austria
+[…]
+[third]organization=Department of Physics, Norwegian University of Science and Technology, addressline=, city=Trondheim, postcode=7491, country=Norway
+[…]
+[fifth]organization=ATLA lasers AS, addressline=Richard Birkelands vei 2B, city=Trondheim, postcode=7034, country=Norway
 1 Introduction
-The unprecedented progress in the field of mesoscopic physics in its general context, which comprises control of macro-quantum states like Bose-Einstein condensate (BEC) [1, 2, 3, 4] and quantum supersolids [5], plasma and turbulence [6, 7, 8], attosecond [9], relativistic and high-energy physics [10, 11, 12, 13], metaphorical modeling and quantum computations [14, 15, 16, 17], physics of ultra-strongly coupled light-matter systems [18], nano-processing [19, 20, 21], and other ranges of science, medicine, and technology, became feasible due to the growth understanding of strongly nonlinear phenomena in far-from-equilibrium systems [22, 23].
+The unprecedented progress in the field of mesoscopic physics in its general context, which comprises control of macro-quantum states like Bose-Einstein condensate (BEC) Kevrekidis et al.
+(2008); Luo et al.
+(2021); Cazalilla et al.
+(2011); Carretero-González et al.
+(2008) and quantum supersolids Boninsegni and Prokof’ev (2012), plasma and turbulence Robinson (1997); Dyachenko et al.
+(1989); Kuznetsov et al.
+(1986), attosecond Krausz and Ivanov (2009), relativistic and high-energy physics Mourou et al.
+(2006); England et al.
+(2014); Südmeyer et al.
+(2008); Gallerati et al.
+(2022), metaphorical modeling and quantum computations Faccio (2012); Kalashnikov and Wabnitz (2021); Faccio et al.
+(2010); Averin et al.
+(2012), physics of ultra-strongly coupled light-matter systems Forn-Díaz et al.
+(2019), nano-processing Gattass and Mazur (2008); Andrews et al.
+(2010); Vorobyev and Guo (2013), and other ranges of science, medicine, and technology, became feasible due to the growth understanding of strongly nonlinear phenomena in far-from-equilibrium systems Cross and Hohenberg (1993); Fu et al.
+(2018).
 The coherent, self-organized, and well-localized patterns, namely solitons, play a crucial role in most of these phenomena.
-The class of such structures, dissipative solitons (DSs), is particularly important because DS develops in open systems, where energy exchange with an environment defines its integrity, stability, and coherence [24, 25].
-In photonics, DSs are a road to generating ultrafast, robust, and energy-scalable pulses in mode-locked lasers [26, 27] bringing a high-field physics on tabletops of a mid-level university lab [12, 28].
-The modern tendencies in generating and exploring laser DSs are based on a controllable enhancement of self-organizing effects induced by nonlinearities and affected both temporal and spatial degrees of freedom [23].
-A partial realization of this concept is a distributed Kerr-lens mode-locking (DKLM) [29, 30, 31].
-In a solid-state laser, this technique uses a nonlinear medium that affects a laser beam via self-focusing, thereby changing an effective gain in a laser resonator [32].
+The class of such structures, dissipative solitons (DSs), is particularly important because DS develops in open systems, where energy exchange with an environment defines its integrity, stability, and coherence Purwins et al.
+(2010); Akhmediev and Ankiewicz (2008).
+In photonics, DSs are a road to generating ultrafast, robust, and energy-scalable pulses in mode-locked lasers Brabec and Krausz (2000); Grelu and Akhmediev (2012) bringing a high-field physics on tabletops of a typical university lab Südmeyer et al.
+(2008); Brons (2017).
+The modern tendencies in generating and exploring laser DSs are based on a controllable enhancement of self-organizing effects induced by nonlinearities and affected by both temporal and spatial degrees of freedom Fu et al.
+(2018).
+A partial realization of this concept is a distributed Kerr-lens mode-locking (DKLM) Sorokina et al.
+(1996); Zhang et al.
+(2022); Demesh et al.
+(2023).
+In a solid-state laser, this technique uses a nonlinear medium that affects a laser beam via self-focusing, thereby changing an effective gain in a laser resonator Brabec et al.
+(1992).
+It was demonstrated that one can even completely decouple the role of the nonlinear medium and the gain medium Sorokina et al.
+(1996).
+The latter opens up the way to power scaling of ultra-short pulsed lasers, especially in waveguide configuration.
 The transversal spatial structure of a field in this type of device is close to a fundamental mode of laser oscillator and, thus, trivial, so only longitudinal modes need synchronization.
-A remarkable breakthrough in energy scalability of such oscillators was demonstrated in both anomalous and normal group-delay dispersion regimes (ADR and NDR, respectively) [33, 34, 30].
-It was found that such broad-range scalability could be explained by energy out/in-flows induced by the nontrivial phase structure of DS [27, 35], and the underlying theory of the DS resonance (DSR) [36] were developed.
-Using DS in the NDR fiber lasers [37] exploits the enhanced and well-controllable nonlinearities.
-That bridges fiber and solid-state ultrafast laser photonics [23].
-The keystone is utilizing a nonlinear propagation of many interacting spatial modes (e.g., in the so-called multimode fibers, MMF) [38, 39].
-Such a nonlinear spatial mode coupling could be caused by nonlinear refraction (or attractive boson interaction in BEC [4]) characterized by the coefficient $n_{2}$.
-A “confining potential” defined by spatial-dependent refractive index $n(x,y)$ (“graded refractive index fibers,” or GRIN, nonlinear lattices [40, 41], or laser-induced confinement in BEC) enhances the relaxation of the higher-order modes to a ground-state, i.e., lowest-order spatial mode.
-In this work, we propose utilizing a nonlinear mode condensation for DKML aimed at a spatio-temporal mode-locking (STML) [42, 43].
-Such an approach utilizes a dissipative (i.e., $\Im[n(x,y)]\neq 0$, where $n$ is, e.g., a complex index of refraction in optics) trapping potential [44] for a spatiotemporal DS (STDS) without involving other nonlinear dissipative processes.
+A remarkable breakthrough in energy scalability of such oscillators was demonstrated in both anomalous and normal group-delay dispersion regimes (ADR and NDR, respectively) Baer et al.
+(2012); Pronin et al.
+(2012); Zhang et al.
+(2022).
+It was found that such broad-range scalability could be explained by energy out/in-flows induced by the nontrivial phase structure of DS Grelu and Akhmediev (2012); Kalashnikov (2012), and the underlying theory of the DS resonance (DSR) Chang et al.
+(2008) was developed.
+Using DS in the NDR fiber lasers Wise et al.
+(2008) exploits the enhanced and well-controllable nonlinearities.
+That bridges fiber and solid-state ultrafast laser photonics Fu et al.
+(2018).
+The keystone is utilizing a nonlinear propagation of many interacting spatial modes (e.g., in the so-called multimode fibers, MMF) Mondal et al.
+(2020); Piccardo et al.
+(2021).
+Such a nonlinear spatial mode coupling could be caused by nonlinear refraction (or attractive boson interaction in BEC Carretero-González et al.
+(2008)) characterized by the coefficient $n_{2}$.
+A “confining potential” defined by spatial-dependent refractive index $n(x,y)$ (“graded refractive index fibers,” or GRIN, nonlinear lattices Matuszewski et al.
+(2006); Kartashov et al.
+(2011), or laser-induced confinement in BEC) enhances the relaxation of the higher-order modes to a ground-state, i.e., lowest-order spatial mode.
+In this work, we propose utilizing a nonlinear mode condensation for DKML aimed at a spatio-temporal mode-locking (STML) Wright et al. (2017, 2020).
+Such an approach utilizes a dissipative (i.e., $\Im[n(x,y)]\neq 0$, where $n$ is, e.g., a complex index of refraction in optics) trapping potential Siegman (2003) for a spatiotemporal DS (STDS) without involving other nonlinear dissipative processes.
 The 2D dissipative potential, unlike that in Refs.
-[45, 46] transversal $(x,y)$-confinement induced by a complex $n(x,y)$ can be associated with a so-called “cigar-type” trapping potential in a weakly dissipative BEC or GRIN (see Table 1) [15].
+Sun et al.
+(2023); Li et al.
+(2022) transversal $(x,y)$-confinement induced by a complex $n(x,y)$ can be associated with a so-called “cigar-type” trapping potential in a weakly dissipative BEC or GRIN (see Table 1) Kalashnikov and Wabnitz (2021).
 Then, an evolutionary (“slow”) coordinate $T$ is associated with time for BEC or propagation distance $Z$ in a fiber or cavity round-trip in a laser.
-A further step is to introduce “longitudinal” (“slow” time $T$-dependent, or propagation length $Z$-dependent) modulation of $\Re[n(x,y,T)]$ (or $\Re[n(x,y,Z])$, which is one of the management techniques stabilizing an ST soliton in BEC [47].
+A further step is to introduce “longitudinal” (“slow”) $T$-dependent (or propagation length $Z$-dependent) modulation of $\Re[n(x,y,T)]$ (or $\Re[n(x,y,Z])$, which is one of the management techniques stabilizing an ST soliton in BEC Kengne et al.
+(2021).
 In a mode-locked laser, it is also possible to identify a “fast” coordinate (local time $t$) associated with the frame co-moving with a soliton.
 The dependence of trapping potential on that coordinate allows a 3D confinement corresponding to a “pancake-like” potential in BEC.
-In a laser, this corresponds to an active mode-locking driven by an external synchronous phase modulation[48, 49, 50], which varies as $\Re[\Delta n(t)]\propto\cos(2\pi t/T_{R})$, where $T_{R}$ is a cavity round-trip period.
+In a laser, this corresponds to an active mode-locking driven by an external synchronous phase modulation Smith et al.
+(1993); Wabnitz (1993); Chang et al.
+(2009), which varies as $\Re[\Delta n(t)]\propto\cos(2\pi t/T_{R})$, where $T_{R}$ is a cavity round-trip period.
 This modulation is slow (pulse duration $\tau_{p}$ is much shorter than $T_{R}$) and can be described by $\Re[\Delta n(t)]\propto t^{2}+\mathrm{H.O.T.}$.
-It was found that such a modulation stabilizes a driven cavity DS [51, 52] and STDS [53].
-Here, we intend to exploit a concept of 3D dissipative confinement to develop a spatiotemporal mode-locking mechanism to generate stable and energy-scalable STDS (“light bullet”).
-This technique can be considered as a road to forming a stable mass-scalable STDS in BEC and pattern formation in dissipative turbulent systems, which could be modeled by the dissipative version of the Gross-Pitaevskii equation (GPE) [4, 54].
-In contrast to [55, 53], a transversely graded dissipation (i.e., $\Im[n(x,y)]\neq 0$) will be supplemented by dissipative trapping along $t$-coordinate ($\Im[n(t)\neq 0$), which is a result of periodical loss modulation (or synchronous pumping) synchronized with the laser cavity round-trip.
+It was found that such a modulation stabilizes a driven cavity DS Leo et al.
+(2010); Englebert et al.
+(2021) and STDS Kalashnikov and Wabnitz (2022) [^1] We note that the driving (modulating) “force” can differ: synchronous modulation of complex nonlinear refraction index (i.e., phase/loss in an oscillator), synchronous pumping, and applying an external coherent pulse.
+One must comprehend the inherent physical differences between them..
+Here, we intend to exploit a concept of 3D dissipative confinement to develop an STML mechanism to generate stable and energy-scalable STDS (“light bullet”).
+This technique can be considered as a road to forming a stable mass-scalable STDS in BEC and pattern formation in dissipative turbulent systems, which could be modeled by the dissipative version of the Gross-Pitaevskii equation (GPE) Carretero-González et al.
+(2008); Zhu et al.
+(2022).
+In contrast to Kalashnikov and Wabnitz (2020, 2022), a transversely graded dissipation (i.e., $\Im[n(x,y)]\neq 0$) will be supplemented by dissipative trapping along $t$-coordinate ($\Im[n(t)]\neq 0$), which is a result of periodic loss modulation (or synchronous pumping) synchronized with the laser cavity round-trip.
 For BEC, this corresponds to a “pancake-like” dissipative confining potential.
-Table 1: Correspondences between photonics and BEC [15]
-Laser
-BEC
-Propagation coordinate $Z$
-$\leftrightarrow$
-time $T$
-Pulse local time $t$
-$\leftrightarrow$
-third spatial coordinate $z$
-Diffraction + anomalous GDD
-$\leftrightarrow$
-boson kinetic energy
-Kerr-nonlinearity
-$\leftrightarrow$
-boson attractive colliding potential
-GRIN
-$\leftrightarrow$
-cigar-like ($x$, $y$) trapping potential
-Non-selective loss
-$\leftrightarrow$
-homogeneous BEC dissipation
-Graded gain + local time modulation
-$\leftrightarrow$
-pancake ($x,y,z$)-shaped inflow of bosons
-Spectral dissipation
-$\Leftrightarrow$
-“kinetic cooling”
+Table 1: Correspondences between photonics and BEC Kalashnikov and Wabnitz (2021)
+Laser BEC
+Propagation coordinate $Z$leftrightarrow time $T$
+Pulse local time $t$leftrightarrow third spatial coordinate $z$
+Diffraction + anomalous GDD $\leftrightarrow$ boson kinetic energy
+Kerr-nonlinearity $\leftrightarrow$ boson attractive colliding potential
+GRIN $\leftrightarrow$ cigar-like ($x$, $y$) trapping potential
+Non-selective loss $\leftrightarrow$ homogeneous BEC dissipation
+Graded gain + local time modulation $\leftrightarrow$ pancake ($x,y,z$)-shaped inflow of bosons
+Spectral dissipation $\leftrightarrow$ “kinetic cooling”
 2 A “soft-aperture” DKLM laser
@@
 The $t-$coordinate corresponds to a “local time” (in a coordinate frame co-moving with an STDS) in photonics or to a longitudinal space coordinate for BEC (see Fig.
-1 and Table 1) [15].
-Figure 1: Multi-scale scheme with a local-time coordinate $t$ synchronized with an STDS group velocity and a global evaluation coordinate $Z$ (see Table 1).
-The gain (filled by a red) is profiled along the radial coordinate $r$ and the local time $t$.
+1 and Table 1) Kalashnikov and Wabnitz (2021).
+Figure 1: Top: Multi-scale scheme with a local-time coordinate $t$ synchronized with an STDS group-velocity and a global evaluation coordinate $Z$ (see Table 1).
+The gain (filled by a red) is profiled along the radial coordinate $\sqrt{x^{2}+y^{2}}$ and the local time $t$.
 The last is synchronized with a laser round-trip period.
-Such pump beam acts as a “soft aperture” guiding the laser beam and controlling (“cleaning”) its mode structure so that a laser field $a(z,x,y,t)$ tends to condensate into a lowest-order mode.
+Bottom: Synchronously pumped waveguide oscillator with a transversely graded refractive index.
+An oscillator can be linear (e.g., solid-state Cr2+:ZnS) or closed (an all-fiber laser).
+A pump pulse could be, for instance, an electrical pulse driving the transversely graded gain coefficient, an optical pump pulse, or an external driving pulse.
+Such pump beam acts as a “soft aperture” guiding the laser beam and controlling its mode structure so that the laser field $a(z,x,y,t)$ tends to condensate into the lowest-order mode.
+This mechanism is akin to DKLM used to generate energy-scalable DS in solid-state lasers when a beam self-focusing enhances a spatial overlapping between pump and generation modes Zhang et al.
+(2022).
+Also, it resembles a beam-by-beam control reported in Ferraro et al.
+(2023).
+However, at variance with the latter, the beam-by-beam coupling in our case is provided not only by Kerr-nonlinearity but also spatially-profiled dissipation.
 Simultaneously, $t-$modulation provides confinement in the time domain.
-Thus, this scheme corresponds to a three-dimensional (“pancake-like”) dissipative confinement for a BEC or a hybrid mode-locking under synchronous pumping or an active assisting mode locking by synchronous loss modulation [56, 57, 50, 52].
+Thus, this scheme corresponds to a three-dimensional (“pancake-like”) dissipative confinement for a BEC or a hybrid mode-locking under synchronous pumping or an active assisting mode locking by synchronous loss modulation Skarka et al.
+(2010); Lobanov et al.
+(2012); Chang et al.
+(2009); Englebert et al.
+(2021).
+Also, $G(x,y,t)$ could be treated as one presenting a first-order expansion of a Gaussian external driving pulse.
 For modeling a (3D+1)-dimensional trapped STDS, we use a distributed model describing the evolution of the $(z,x,y,t)$-dependent field $a$ propagating along the $z$-axis (see Table 1 for comparison with BEC) of an active waveguide with the linear and nonlinear refractive indexes $n(x,y)$ and $n_{2}$, respectively.
 Under the action of $(x,y,t)$-graded pump (Fig.
-1), the evolution can be described by a sort of the driven nonlinear Helmholtz equation:
+1), the evolution can be described by a variation of the driven nonlinear Helmholtz equation:
 $-\left[\frac{1}{2k_{0}n_{0}}\left(\Delta_{x,y}-\frac{\partial^{2}}{\partial z^{2}}\right)+(\frac{\beta_{2}}{2}-i\tau)\frac{\partial^{2}}{\partial t^{2}}\right]a=$ (1)
 $\left[\frac{k_{0}}{2n_{0}}\left(n(x,y)^{2}-n_{0}^{2})\right)+i\left(L-G(x,y,t)\right)+k_{0}n_{2}\left|a\right|^{2}+i\frac{\partial}{\partial z}\right]a,$
-where $\Delta_{x,y}$ is a Laplace operator in the Cartesian coordinates, $n_{0}$ is a waveguide cladding refractive index, $k_{0}$ is a free-space wave number, and the second-order derivative over $z$ contributes only for large beam numerical apertures NA [58, 59].
+where $\Delta_{x,y}$ is a Laplace operator in the Cartesian coordinates, $n_{0}$ is a waveguide cladding refractive index, $k_{0}$ is a free-space wave number, and the second-order derivative over $z$ contributes only for large beam numerical apertures NA Akhmanov et al.
+(1992); Demesh et al.
+(2023).
 The terms in Eq.
-1 define: (I) – diffraction, (II) – group-delay dispersion with the coefficient $\beta_{2}$, (III) – spectral dissipation with the squared inverse bandwidth $\tau$, (IV) – trapping potential induced by GRIN, (V) – $(x,y,t)$- graded saturated gain with the coefficient $G$ and linear net-loss $L$, and (VI) – Kerr-nonlinearity.
+(1) define: (I) – diffraction, (II) – group-delay dispersion with the coefficient $\beta_{2}$, (III) – spectral dissipation with the squared inverse bandwidth $\tau$, (IV) – trapping potential induced by GRIN, (V) – $(x,y,t)$- graded saturated gain with the coefficient $G$ (see footnote 1) and linear net-loss $L$, and (VI) – Kerr-nonlinearity.
 One has to emphasize the decisive contribution of a saturable and transversely profiled gain $G(x,y,t)$ in Eq.
@@
 (1).
-That leads to the nonlinear driven Schrödinger, or the Lugiato-Lefever equation (LLE), which is a test bed model for the study of solitons and non-equilibrium patterns in driven Kerr resonators, micro-cavity and fiber lasers, VCSEL, etc.
-[60, 61, 62, 63].
-The rescaling $\psi=a\exp(-iV_{0}z)$ ($V_{0}=(k_{0}/2n_{0})(n_{1}^{2}-n_{0}^{2})$, and $n_{1}$ is a waveguide core refractive index) leads to the dimensionless mean-field LLE [64, 65, 66]:
+That leads to the nonlinear driven Schrödinger or Gross-Pitaevskii equation [^2] The terminology depends on the physical treatment of $G(x,y,t)$.
+It could be a gain coefficient created by synchronous pumping, active gain modulation, or a driving external pulse.
+In the last case, the master equation resembles the Lugiato-Lefever equation (LLE), which is a test bed model for the study of solitons and non-equilibrium patterns in driven Kerr resonators, micro-cavity and fiber lasers, VCSEL, etc.
+Lugiato et al.
+(2015); Castelli et al.
+(2017); Lugiato et al.
+(2018); Coen and Erkintalo (2016).
+The rescaling $\psi=a\exp(-iV_{0}z)$ ($V_{0}=(k_{0}/2n_{0})(n_{1}^{2}-n_{0}^{2})$, and $n_{1}$ is a waveguide core refractive index) leads to the dimensionless mean-field equation:
 $i\frac{\partial\psi\left(Z,X,Y,t^{\prime}\right)}{\partial Z}=\left\{-\frac{1}{2}\left[\frac{\partial^{2}}{\partial X^{2}}+\frac{\partial^{2}}{\partial Y^{2}}+\frac{\partial^{2}}{\partial t^{\prime 2}}-2i\tau\frac{\partial^{2}}{\partial t^{\prime 2}}\right]\right\}\psi+$
-$\left\{(X^{2}+Y^{2})-\left|\psi\right|^{2}-i\left[\Lambda+\kappa(X^{2}+Y^{2})+\omega t^{\prime 2}\right]\right\}\psi,$ (2)
-for the normalizations shown in Table 2 [64, 67].
+$\left\{\frac{1}{2}(X^{2}+Y^{2})-\left|\psi\right|^{2}-i\left[\Lambda+\kappa(X^{2}+Y^{2})+\omega t^{\prime 2}\right]\right\}\psi,$ (2)
+for the normalizations shown in Table 2 Lugiato and Lefever (1987); Raghavan and Agrawal (2000); Haelterman et al. (1992).
 The dimensional extension by the “fast-time” (or “local time”) coordinate $t^{\prime}$ allows describing the pulse dynamics in the co-moving frame $(Z,X,Y,t^{\prime})$.
@@
 The normalized version (2) assumes an anomalous net-GDD.
-In this work, we assume the parabolic-like transverse profile of the gain beam with the size.
-Such transversely confined nonlinear gain acts as a “soft aperture” with an effective size $\approx w_{p}\sqrt{G_{0}-\Lambda}$ (Table 2).
-Our model introduces a gain modulation along the $t-$axis described in a parabolic approximation by the modulation coefficient $\omega$ [50, 57].
-The last resembles the mechanism of the so-called active mode-locking (AM) [68].
-In contrast to [53], the proposed method does not use additional phase modulation but only a pumping pulse train synchronized with a resonator period.
+In this work, we assume the parabolic-like transverse profile of the gain beam.
+Such transversely confined gain acts as a “soft aperture” with an effective size $\approx w_{p}\sqrt{G_{0}-\Lambda}$ (Table 2).
+Our model introduces a gain modulation along the $t-$axis described in a parabolic approximation by the modulation coefficient $\omega$ Chang et al.
+(2009); Lobanov et al.
+(2012).
+The last resembles the mechanism of the so-called active mode-locking Herrmann and Wilhelmi (1987).
+In contrast to Kalashnikov and Wabnitz (2022), the proposed method does not use additional phase modulation but only a pumping pulse train synchronized with a resonator period.
 For simplicity, we manipulate with a graded saturated gain $G(X,Y,t)=G(0,0,0)\times\left[1-\kappa(X^{2}+Y^{2})-\omega t^{\prime 2}\right]$, where $G(0,0,0)$ or $G_{0}$ (see Table 2) corresponds to a saturated gain along a waveguide axis at $t^{\prime}=0$, and $\kappa=G_{0}w_{p}^{-2}$, $\omega=G_{0}\mathscr{T}^{-2}$ ($\mathscr{T}$ is a “pump pulse duration”).
-Such an approximation is valid under the condition of quasi-steady state operation of a laser in the vicinity of the lasing threshold, where the saturated net-loss coefficient $\Lambda^{\prime}=L-G_{0}$ is close to zero [55].
+Such an approximation is valid under the condition of quasi-steady state operation of a laser in the vicinity of the lasing threshold, where the saturated net-loss coefficient $\Lambda^{\prime}=L-G_{0}$ is negative and close to zero Kalashnikov and Wabnitz (2020).
 Another essential factor in Eqs.
@@
 This parameter is inversely proportional to a squared spectral width of the gain band or spectral filter (Table 2).
-Thus, one may treat the resulting equation (2) as a dissipative extension of the driven Gross-Pitaevskii equation, which is a well-known tool for analyzing trapped BEC (Table 1) [4].
-Table 2: Normalizations for the LLE (2)
-Value
-Normalizations and definitions
-$(X,Y)$
-transverse coordinates, $(x,y)/w_{T}$
-$Z$
-propagation distance, $z/\zeta$
-$t^{\prime}$
-local time, $t/\sqrt{|\beta_{2}|\zeta}$
-$\zeta$
-propagation scale, $k_{0}n_{o}w_{T}^{2}$
-$w_{T}$
-transverse scale, $4]{w_{0}^{2}/k_{0}^{2}n_{0}\delta}$
-$\delta$
-GRIN contrast, $n_{1}-n_{0}$
-$\left(k_{0}/2n_{0}\right)\left(n_{1}^{2}-n(x,y)^{2})\right)$
-GRIN confining potential,  $\simeq k_{0}\left(n_{1}-n(r)\right)$
-$\left|\psi\right|^{2}$
-intensity, $k_{0}n_{2}\zeta\left|a\right|^{2}$
-$\Lambda^{\prime}$
-net-loss, $\Lambda-G_{0}$
-$\Lambda$
-averaged loss, $LL_{w}/\zeta$
-$\kappa$
-spatial confinement parameter, $G_{0}/w_{p}^{2}$
-$\omega$
-temporal confinement parameter, $G_{0}/\mathscr{T}^{2}$
-$w_{p}\sqrt{G_{0}-\Lambda}$
-soft aperture size
-$\tau$
-squared inverse spectral filter, bandwidth normalized to $|\beta_{2}|\zeta$
-$k_{0}$
-wave-number, $2\pi/\lambda$
-$\lambda$
-central wavelength
-$n_{1}$
-GRIN core refractive index
-$n_{0}$
-GRIN cladding refractive index
-$n_{2}$
-nonlinear refractive index
-$L$
-loss coefficient
-$L_{w}$
-waveguide (“lattice”) length
-$G_{0}$ or $G(t=0,x=0,y=0)$
-maximum saturated gain coefficient
-$w_{p}$
-pump beam size
-$\mathscr{T}$
-pump pulse width
+Thus, one may treat the resulting equation (2) as a dissipative extension of the driven Gross-Pitaevskii equation, which is a well-known tool for analyzing trapped BEC (Table 1) Carretero-González et al.
+(2008).
+Table 2: Normalizations for Eq. (2)
+Value Normalizations and definitions
+$(X,Y)$ transverse coordinates, $(x,y)/w_{T}$
+$Z$ propagation distance, $z/\zeta$
+$t^{\prime}$ local time, $t/\sqrt{|\beta_{2}|\zeta}$
+$\zeta$ propagation scale, $k_{0}n_{o}w_{T}^{2}$
+$w_{T}$ transverse scale, $4]{w_{0}^{2}/k_{0}^{2}n_{0}\delta}$
+$\delta$ GRIN contrast, $n_{1}-n_{0}$
+$\left(k_{0}/2n_{0}\right)\left(n_{1}^{2}-n(x,y)^{2})\right)$ GRIN confining potential,  $\simeq k_{0}\left(n_{1}-n(r)\right)$
+$\left|\psi\right|^{2}$ intensity, $k_{0}n_{2}\zeta\left|a\right|^{2}$
+$\Lambda^{\prime}$ net-loss, $\Lambda-G_{0}$
+$\Lambda$ averaged loss, $LL_{w}/\zeta$
+$\kappa$ spatial confinement parameter, $G_{0}/w_{p}^{2}$
+$\omega$ temporal confinement parameter, $G_{0}/\mathscr{T}^{2}$
+$w_{p}\sqrt{G_{0}-\Lambda}$ soft aperture size
+$\tau$ squared inverse spectral filter, bandwidth normalized to $|\beta_{2}|\zeta$
+$k_{0}$ wave-number, $2\pi/\lambda$
+$\lambda$ central wavelength
+$n_{1}$ GRIN core refractive index
+$n_{0}$ GRIN cladding refractive index
+$n_{2}$ nonlinear refractive index
+$L$ loss coefficient
+$L_{w}$ waveguide (“lattice”) length
+$G_{0}$ or $G(t=0,x=0,y=0)$ maximum saturated gain coefficient
+$w_{p}$ pump beam size
+$\mathscr{T}$ pump pulse width
 2.2 Variational approximation
@@
 (2) is based on analytical and numerical approaches.
-The first uses the well-known variational approximation (VA) [55].
-The generating Lagrangian for a non-dissipative part of (2) is
-$\mathscr{L}=\frac{i}{2}\left(\psi^{*}\partial_{Z}\psi-\psi\partial_{Z}\psi^{*}\right)+\frac{1}{2}\left(\left|\partial_{X}\psi\right|^{2}+\left|\partial_{Y}\psi\right|^{2}+\left|\partial_{t^{\prime}}\psi\right|^{2}\right)-$
-$\left|\psi\right|^{4}+\left(X^{2}+Y^{2}\right)\left|\psi\right|^{2}.$ (3)
-The dissipative factors can be described in the form of a “force” source term
+The first uses the well-known variational approximation (VA) Karlsson et al.
+(1992); Yu et al.
+(1995); Malomed (2002); Kalashnikov and Wabnitz (2020); Parra-Rivas et al.
+(2023).
+The generating Lagrangian density for a non-dissipative part of (2) is
+$\mathscr{L}=\frac{1}{2}\left\{i\left(\psi^{*}\partial_{Z}\psi-\psi\partial_{Z}\psi^{*}\right)+\left(\left|\partial_{X}\psi\right|^{2}+\left|\partial_{Y}\psi\right|^{2}+\left|\partial_{t^{\prime}}\psi\right|^{2}\right)\right\}-$
+$\frac{1}{2}\left|\psi\right|^{4}+\frac{1}{2}\left(X^{2}+Y^{2}\right)\left|\psi\right|^{2}.$ (3)
+The dissipative factors can be described in the form of the Rayleigh dissipative function
 $\mathcal{Q}=-i\left[\Lambda^{\prime}+\kappa(X^{2}+Y^{2})+\omega t^{\prime 2}-\tau\frac{\partial^{2}}{\partial t{{}^{\prime 2}}}\right]\psi$ (4)
-in the Euler-Lagrange equation
+in the Euler-Lagrange-Kantorovich equation
 […]
-in agreement with Kantorovitvh’s method (see [55, 15, 53] and references herein).
+in agreement with Kantorovich’s method, which is a generalization of the Rayleigh-Ritz method Kantorovich and Krylov (1958); Chávez Cerda et al.
+(1998).
+In Eq.
+(5), the variation is made over the $Z-$dependent parameters ${\mathop{\rm f}\nolimits}$ of some ansatz, which will be introduced below.
+The reduced Lagrangian
+[…]
+is evaluated after substituting an ansatz in $L$ and $Q$.
 It is convenient to assume the axial symmetry, which corresponds to the symmetry of a cylindrical waveguide, and transit to a radial coordinate $R=\sqrt{X^{2}+Y^{2}}$.
-The next assumption is a Gaussian-sech approximation for an STDS profile:
+The next assumption is a Gaussian-sech ansatz for Eqs.
+(5,6) Yu et al.
+(1995):
 $\psi(Z,R,t^{\prime})=\alpha(Z)\operatorname{sech}\left[\frac{t^{\prime}}{\Upsilon(Z)}\right]\times$
@@
 (5) are: $\alpha$ amplitude, $\phi$ phase, $\chi$ and $\theta$ are the temporal and spatial chirps, respectively; $\rho$ and $\Upsilon$ are the beam size and STDS duration, respectively.
-After some algebraic manipulations, one can obtain the expressions for the steady-state (i.e., $Z-$independent) STDS parameters (the phase $\phi$ is irrelevant in our context):
-$\rho^{2}=\frac{1}{384\pi^{2}\kappa\tau\Upsilon^{2}}\times$ (7)
-$\biggl{[}15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}-15\right)-$
-$64\tau\left(15\tau+2\pi^{2}\left(\tau+3\Lambda^{\prime}\Upsilon^{2}\right)\right)\biggr{]},$
-$\chi=\frac{\sqrt{\Upsilon^{4}\left(64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225\right)}-15\Upsilon^{4}}{16\pi^{2}\tau\Upsilon^{4}},$ (8)
-$\alpha_{0}^{2}=\frac{-6\kappa^{2}\rho^{8}-6\rho^{4}+6}{2\rho^{2}},$
-$\theta=-\frac{\kappa\rho^{2}}{2}.$
+We use this ansatz as an approximation for a pulse with the near-Gaussian transverse spatial profile, i.e., the lowest-order spatial mode in the parabolic potential of Eq.
+(3), and a sech-shaped pulse with the quadratic phase for a temporal profile.
+The last was widely used for VA.
+Other ansatzs with different temporal phase profiles were also considered (e.g., see Malomed (2002).
+In our case, the temporal shape of ansatz is conditioned by the following.
+This shape corresponds to a dissipative soliton solution $\operatorname{sech}(t^{\prime}/\Upsilon)^{1-i\psi}$ of Eq.
+(2) without a confining potential Maimistov and Skliarov (1987) in a limit when the exact phase profile $-\psi\ln{(\operatorname{sech}(t^{\prime}/\Upsilon)})\approx\psi t^{\prime 2}/2\Upsilon^{2}$ (here, $\psi=2\Upsilon^{2}\chi$ is a dimensionless chirp).
+After some algebraic manipulations (see Appendix A and Kalashnikov (2024)), one can obtain the expressions for the steady-state (i.e., $Z-$independent) STDS parameters (there are two solutions for the chirp parameter $\chi$, but one of them is nonphysical):
+$\rho_{0}^{2}=\frac{1}{384\pi^{2}\kappa\tau\Upsilon_{0}^{2}}\times$
+$\biggl{[}15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225}-15\right)-$
+$64\tau\left(15\tau+2\pi^{2}\left(\tau+3\Lambda^{\prime}\Upsilon_{0}^{2}\right)\right)\biggr{]},$
+$\chi_{0}=\frac{\sqrt{\Upsilon_{0}^{4}\left(64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225\right)}-15\Upsilon_{0}^{4}}{16\pi^{2}\tau\Upsilon_{0}^{4}},$ (8)
+$\alpha_{0}^{2}=\frac{-6\kappa^{2}\rho_{0}^{8}-6\rho_{0}^{4}+6}{2\rho^{2}},$
+$\theta_{0}=-\frac{\kappa\rho_{0}^{2}}{2}.$
 The remaining equation for the STDS duration $\Upsilon$
-$144\left(\pi^{2}-25\right)\tau^{2}+\left(16\left(3+\pi^{2}\right)\tau^{2}+45\right)\times$ (9)
-$\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}+96\pi^{4}\Upsilon^{4}\omega+$
-$+\frac{2^{13}3^{3}\pi^{4}\kappa\tau^{3}\Upsilon^{4}\left(-\frac{A^{4}}{3\times 2^{15}\pi^{8}\kappa^{2}\tau^{4}\Upsilon^{8}}-\frac{3A^{2}}{2\pi^{4}\kappa^{2}\tau^{2}\Upsilon^{4}}+1\right)}{A}=675,$ (10)
-$A=64\tau\left(2\pi^{2}\left(3\Lambda^{\prime}\Upsilon^{2}+\tau\right)+15\tau\right)-$
-$15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}-15\right)$
-must be solved numerically.
+$144\left(\pi^{2}-25\right)\tau^{2}+\left(16\left(3+\pi^{2}\right)\tau^{2}+45\right)\times$
+$\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225}+96\pi^{4}\Upsilon_{0}^{4}\omega\tau+$
+$+\frac{2^{13}3^{3}\pi^{4}\kappa\tau^{3}\Upsilon_{0}^{4}\left(-\frac{A^{4}}{3\times 2^{15}\pi^{8}\kappa^{2}\tau^{4}\Upsilon_{0}^{8}}-\frac{3A^{2}}{2\pi^{4}\kappa^{2}\tau^{2}\Upsilon_{0}^{4}}+1\right)}{A}=675,$ (9)
+$A=64\tau\left(2\pi^{2}\left(3\Lambda^{\prime}\Upsilon_{0}^{2}+\tau\right)+15\tau\right)-$
+$15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225}-15\right)$
+must be solved numerically. Below, we will omit the 00-subscribe for steady-state (fixed-point) solutions characterizing STDS.
 2.3 STDS parameters
 Figs.
-2-5 show the STDS width $\Upsilon$ and peak intensity $\alpha_{0}^{2}$ obtained from Eqs.
-(7-9).
-$\Upsilon$ grows, and $\alpha_{0}^{2}$ decreases with $\kappa$ when a spectral dissipation $\tau$ is sufficiently low (Figs.
-2, 3).
-Such a tendency corresponds to that of a DKLM regime with $\omega=0$ and manifests the growth of spatial mode loss with an aperture squeezing [55].
-In the vicinity of a “threshold” ($\Lambda^{\prime}\to$0), the additional longitudinal (3D) trapping defined by $\omega$ noticeably squeezes an STDS with the simultaneous peak power growth (Fig.
-3).
-The additional solution branch in the region of small $\kappa$ (Fig.
-2) disappears with $\Lambda^{\prime}\to$0.
-In general, the STDS intensity growth and its width shortening with $\omega$ could result from an appearance of the external confinement along the $t-$ coordinate.
-The spectral dissipation increases with $\tau$.
-That leads to an STDS collapse for $\kappa\to|\Lambda^{\prime}|$ when the net-loss coefficient $\Lambda^{\prime}$ is more significant in comparison with Fig.
-3.
-Such a collapse is asymmetrical in space and time.
-That is, STDS spreads along $t$ and squeezes on $R$.
-A longitudinal mode selection growth with $\tau$ (i.e., STDS spreading in the time domain) parallels a transverse mode selection (STDS spatial squeezing) in the spatial domain.
-The solutions bifurcate close to the threshold $\Lambda^{\prime}$ (Fig.
-5).
-Such behavior leads to the thought of stability loss here.
-These oddities require numerical simulations of STDS dynamics presented below.
-Figure 2: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters.
+2-4 show the STDS width $\Upsilon$ and peak intensity $\alpha_{0}^{2}$ obtained from Eqs.
+(8, 9).
+$\Upsilon$ grows, and $\alpha_{0}^{2}$ decreases with $\kappa$.
+Such a tendency corresponds to that of a DKLM regime with $\omega=0$ and manifests the growth of spatial mode loss with a soft aperture squeezing $\kappa=G_{0}w_{p}^{-2}$ Kalashnikov and Wabnitz (2020).
+The nonzero $\omega$ contributes to this behavior and squeezes the $\kappa$-range where the STDS exists.
+It could be explained by an additional loss due to confinement in the $t$-domain.
+Insets in Fig.
+4 demonstrate the pulse spatio-temporal and spectral profiles corresponding to the ansatz (7) (see Appendix B).
+The solutions for $\omega\neq$0 bifurcate close to the threshold $\Lambda^{\prime}\to$0 (Figs.
+3,4).
+The new solutions with the opposite sign of the chirp $\chi$ appear.
+The dependencies of $\Upsilon$ and $\alpha^{2}$ on $\kappa$ for these new solutions are opposite to those for $\omega=$0.
+Since the $\tau$-growth increases a spectral loss, it causes the $\alpha^{2}$-decrease and $\Upsilon$-rise (compare Figs.
+3 and 4).
+Figure 2: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ (b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters.
 $\Lambda^{\prime}=$-0.01, $\tau=$0.01.
-Figure 3: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters.
+The unstable solutions are shown by dashed curves.
+A bold magenta curve shows the stable solutions for $\omega=10^{-5}$.
+Figure 3: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters.
 $\Lambda^{\prime}=-$0.001, $\tau=$0.01.
-Figure 4: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters.
-$\Lambda^{\prime}=-$0.01, $\tau=$0.1.
+Dashed curves correspond to unstable solutions.
+Solutions corresponding to $\omega$=10-4 and 10-3 are nonphysical.
 Figure 4: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ (b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters.
 $\Lambda^{\prime}=-$0.001, $\tau=$0.1.
-2.4 STDS stability
+Insets show the analytical spatio-temporal (a) and spectral (b) profiles of STDS for the parameters marked by the green circles ($\kappa=$0.0006, $\omega=$10-7).
+The blue square marks the parameters $\kappa=$0.0005, $\omega=$10-6 corresponding to Fig.
+7.
+The unstable solutions are shown by dashed curves.
+The white square marks the parameters corresponding to Fig.
+8.
+Solutions for $\omega$=10-4 and 10-3 are nonphysical.
+2.4 STDS stability: linear analysis
+The linear stability analysis is based on the standard technique of the study of the Jacobian eigenvalues (see Appendix C and Kalashnikov (2024)) at the stationary points of (13).
+Their real parts must be negative to provide a local stability of the STDS solutions.
+The analysis demonstrates that the solutions presented in Fig.
+2 are unstable, expecting the confined region of $\kappa\to$0 for a non-zero $\omega$ (thick magenta line).
+The dependencies of the corresponding Jacobian eigenvalues $\lambda$ on $\kappa$ are shown in Fig.
+5.
+When $\Lambda^{\prime}\to$0, the additional unstable solutions appear for the considered non-zero $\omega$ as is shown in Figs.
+3,4.
+But in this case, the first solutions corresponding to the branches with lower $\alpha^{2}$ are stable for $\omega\neq$0 within all diapason of $\kappa$ in which an STDS exists.
+This diapason shrinks with $\omega$ because the $\omega$-growth shortens a gain “window” in the $t^{\prime}$-domain.
+The analysis testifies that, for an unstable solution, there is always an eigenvalue $\lambda$ with $\Re(\lambda)>$0 but $\Im(\lambda)=$0 that does not show evidence of the limiting circles.
+On the other side, the numerical simulation of (13) demonstrates a rapid convergence to a stable solution within its existence region, as is shown in Fig.
+6.
+One must note that an absence of analytical solutions does not mean a lack of localized solutions because the model is confined by the choice of the ansatz (7).
+Therefore, the numerical simulations of Eq.
+(2) are required.
+Figure 5: Real and imaginary parts of the Jacobian eigenvalues $\lambda$ in dependence on $\kappa$ for $\omega$=10-5, $\Lambda^{\prime}=-$0.01, and $\tau=$0.01 Kalashnikov (2024).
+Figure 6: Evolution of $\Upsilon(Z)$ and $\alpha^{2}(z)$ for the parameters marked by points in Fig.
+4.
+The initial values are $\Upsilon(0)=$10, $\alpha(0)=$0.1, $\rho(0)=$1, and $\psi(0)=\theta(0)=\phi(0)=$0.
+2.5 STDS stability: numerical simulations
 We performed large-scale numerical simulations of Eq.
@@
 The characteristic case is illustrated in Fig.
-6, where the numerical STDS profile corresponds to one of the VA solutions from Fig.
+7 for the parameters marked by the blue square in Fig.
 4.
-The red line shows the analytical and numerical solutions shown by the black curve.
-One can see an oscillatory approach from the last to the first, analogous to the damping of excitations in a weakly dissipative BEC [69].
-Interestingly, the peak intensity of the numerical solution is slightly lower than the analytical one.
-The difference in the exact solution from the Gaussian ansatz for VA could explain it.
-[…]
-Besides the stable STDS solutions, the numerical simulations demonstrate three main destabilization scenarios.
-The first is the STDS decay with the $|\Lambda^{\prime}|$-decrease.
-The second is the collapse-like behavior for small $\kappa$, $\omega$, and $\tau$.
-Both demonstrate the importance of dissipative factors (i.e., dissipative confinement and spectral filtering) for STDS stability.
-The third scenario is the DS splitting with possible oscillations between multiple pulse states.
+The red line corresponds to the analytical solution, and the black curve shows the numerical solution for the maximal $|\psi|^{2}$ in dependence on $Z$.
+The ansatz (7) with $\alpha_{0}$=0.46, $\Upsilon$=22, $\rho$=1, $\psi=\theta$=0 was chosen as the initial condition.
+One can see a fast approach of the STDS maximal intensity to the analytical solution.
+Figure 7: Evolution of the STDS peak intensity $|\psi(R=0,t^{\prime}=0,Z)|^{2}$ (black solid curve) and the STDS profile $|\psi(R,t^{\prime},Z=3000)|^{2}$ for the parameters of Fig.
+4 (blue square).
+The red line shows the analytical $\alpha_{0}^{2}$ from Fig.
+4 for $\omega=$10-6, $\kappa=$0.0005.
+Besides the stable STDS solutions, the numerical simulations demonstrate two main destabilization scenarios.
+The first is the STDS decay (Fig.
+8) corresponding to an unstable analytical solution of Fig.
+4,b (white square).
+For this scenario, a peak amplitude decreases exponentially with a slow temporal broadening of STDS.
+Figure 8: Evolution of the STDS peak intensity $|\psi(R=0,t^{\prime}=0,Z)|^{2}$ (black solid curve, logarithmic scale), and the initial and final STDS profiles $|\psi(R,t^{\prime},Z=0,10000)|^{2}$ for the parameters of Fig.
+4 marked by the white square.
+The second scenario is the STDS splitting with the irregular oscillations between multiple pulse states leading to a “rogue wave” dynamics.
 Fig.
-7 shows such regimes.
-The splitting in a local time domain could be explained by weaker confinement in the time domain due to lower $\omega$.
-However, a more precise analysis of the STDS requires considering the gain saturation dynamics.
-Figure 7: Contour plot of the STDS peak intensity $\alpha_{0}^{2}$ at $Z=$40000 (a) and 35000 (b) for $\omega=10^{-5}$ (a) and $2\times 10^{-6}$ (b), $\kappa=$0.005, $\Lambda^{\prime}=$-0.01, and $\tau=$0.1.
+9 shows such a regime, which cannot be described within the above analytical model.
+One may conjecture that the splitting in a local time domain could be explained by broadening the confinement potential with the $|\Lambda^{\prime}|$-growth.
+However, a more precise analysis of such dynamics requires considering the gain saturation.
+Figure 9: Evolution of the maximal $|\psi|^{2}_{max}$ with $Z$ for $\omega=10^{-5}$, $\kappa=$0.005, $\Lambda^{\prime}=$-0.01, and $\tau=$0.1.
+Inset shows the profile of the STDS complex at the final point of the considered $Z$-interval.
 2.6 STDS stability under the saturable gain dynamics
-To take into account the dynamical gain saturation, it is convenient to use the Frantz-Nodvik equation [70]:
+To take into account the dynamic gain saturation, it is convenient to use the Frantz-Nodvik equation Frantz and Nodvik (1963):
 $\frac{\partial G_{0}}{\partial t}=\frac{G_{ss}-G_{0}}{T_{r}}-\frac{G_{0}\left|\psi\right|^{2}}{E_{s}},$ (10)
 where $G_{ss}$ is a small signal gain, $T_{r}$ is a gain relaxation time, $E_{s}=h\nu/S\sigma_{em}$ is a gain saturation energy ($\nu$ is a central gain-band wavelength coinciding with the STDS carrier frequency, $S$ is a beam area, $\sigma_{em}$ is a stimulated emission cross-section).
-Specifically, we base on the characteristics of the Cr:ZnS waveguide laser [71].
-Table 3 presents the corresponding dimensional laser parameters.
-Table 3: Correspondences between photonics and BEC
-Cr:ZnS laser parameters (see Table 1) values
-gain relaxation time $T_{r}$, $\mu s$ 4.9
+Specifically, we base on the characteristics of a Cr:ZnS waveguide laser Sorokin et al.
+(2022).
+Table 3 presents the corresponding dimensional laser parameters ($n_{2}$ value is defined at 1.3 $\mu$m).
+Table 3: Cr:ZnS laser parameters Sorokina et al. (2002); Sorokina (2008)
+parameters values
+radiative lifetime $T_{r}$, $\mu s$ 4.3
 […]
 […]
-central wavelength $\lambda$, $\mu$m 2.3
+central wavelength $\lambda$, $\mu$m 2.35
 $w_{0}$, $\mu$m 40
-$n_{0}$ 2.2629
-$\delta$ 0.0077
-numerical aperture $NA=\sqrt{n_{1}^{2}-n_{0}^{2}}$ $\approx 0.19$
+$n_{0}$ at 2.3 $\mu$m 2.27
+$\delta$ 0.0047
+numerical aperture $NA=\sqrt{n_{1}^{2}-n_{0}^{2}}$ $\approx 0.15$
 $w_{T}$, cm 0.001
@@
 […]
+$\beta_{2}$, fs2/cm 1280
 […]
-The calculations demonstrate a visible effect of the gain saturation on the DS dynamics.
-For a small level of linear loss $\Lambda$, some characteristic profiles of STDS are shown in Fig.
-8.
-As one can see, a sufficiently large $\omega$ stabilizes DS ($a$), but the driving frequency decrease leads to the STDS shape distortion, which is periodically asymmetrical along $t$-coordinate.
-Oscillations accompany these distortions and result in STDS decay.
-STDS cannot be stabilized for the parameters considered for a smaller $\Lambda$ (Fig.
-5).
 […]
+For a chosen level of linear loss $\Lambda$, some characteristic profiles of STDS based on the numerical simulations are shown in Figs.
+10.
+As one can see, an additional scenario of STDS destabilization appears.
+Namely, there are the STDS shape distortion, shift, and splitting, which alternate irregularly.
+Figure 10: The $|\psi|^{2}$-profiles on the different distances $Z=$16000 (a), 29000 (b), and 30000 (c) for $\kappa=$0.002, $\omega=10^{-5}$, $\Lambda$=0.1, $G_{ss}$=1.1 $\Lambda$, and $\tau=$0.1 (see the green circle in Fig.
+11).
+Figure 11: The visualization of the STDS stability region on ($\kappa$$\omega$)-plane under dynamic gain saturation.
+The nodes correspond to the “last” numerical stability points connected “by hand” for clarity.
+The green circle corresponds to the parameters of Fig.
+10.
+$\Lambda$=0.1, $G_{ss}$=1.1 $\Lambda$, and $\tau=$0.1.
 The region of STDS stability is shown in Fig.
-9 for a larger linear loss coefficient $\Lambda$.
+11 for a given linear loss coefficient $\Lambda$.
+The “nodes” in Figure correspond to the “last” stability points obtained numerically and are connected “by hand” for clarity.
 The STDS decay confines this region for larger $\kappa$, so there are no stable pulses for $\kappa\gtrsim 0.006$.
@@
 Such a gain reduction is a direct consequence of the gain saturation.
-For smaller $\kappa$, the main destabilization factor is periodical pulse splitting, which alternates with a single-pulse regime (Figs.
-8 (b-d) and inset in Fig.
-9).
-The asymmetry of the STDS complex appears periodically due to dynamic gain saturation.
+The main destabilization factor for smaller $\kappa$ is irregular pulse splitting, which alternates with a single-pulse regime (Figs.
+10).
+The asymmetry of the STDS complex appears due to dynamic gain saturation (Figs.
+10).
 When $\kappa$ approaches zero, very strong STDS power and width oscillations appear.
-They reach more than two orders of magnitude, so STDS is periodically “reborn”.
-This phenomenon confines the stability region from below on $\kappa$.
-Figure 9: Region of STDS stability on ($\kappa$$\omega$)-plane under the action of dynamical gain saturation.
-$\Lambda=$0.1, and other parameters as in Fig.
-8.
-Inset shows the $|\alpha_{0}|^{2}$ contour plot for $\kappa=5\times 10^{-4}$, $\omega=\times 10^{-6}$.
+They reach more than two orders of magnitude (Figs.
+10 (a) vs.
+(b,c)), so STDS is periodically “reborn”.
+This phenomenon confines the stability region from below on $\kappa$ and above on $\omega$.
+As an additional effect that waits for further exploration is a symmetry breaking.
+It was found Kalashnikov and Wabnitz (2022) that this effect contributes to STDS dynamics in a system with pure real pancake-like plus imaginary cigar-like potentials.
+The STDS shape can be distorted, and the dynamics becomes more complicated.
+However, STDS remains 3D-localized.
+The situation is reversed in our case: pure imaginary pancake-like plus real cigar-like potentials.
+The numerical analysis shows that, unlike Kalashnikov and Wabnitz (2022), there is no STDS splitting in the transverse dimension $R$.
+That is, the transverse perturbations are suppressed, and we may conjecture that the development of the spatial asymmetry would be troubled.
 3 Discussion and Conclusion
-Similar to the mechanism of STDS stabilization by non-dissipative confinement along the longitudinal ($t$) coordinate [53], which corresponds to an additive phase mode-locking in a laser, the dissipative confinement due to synchronous pulse pumping or active amplitude mode-locking leads to a stable STDS, as well.
-Such stabilization is possible at very low values of the modulation “depth” parameter $\omega=G_{0}/\mathscr{T}^{2}\approx\Lambda/\mathscr{T}^{2}\approx 10^{-4}$.
-This means that gain variations along a laser axis could be very small.
-For instance, it means that a pumping pulse, synchronized with the cavity round-trip, could be very “long” compared to the STDS temporal width.
-The contribution of dynamical gain saturation lowers this value to $\approx 10^{-6}$.
-Simultaneously, there is a lower threshold on the value of the saturated net gain $|\Lambda^{\prime}|$.
-In our case, it was $\approx 0.01$.
-Too low net gain could not provide the 3D dissipative confinement and, thereby, STDS stabilization.
-We also found that the most optimal stabilization region is located near $\kappa\approx|\Lambda^{\prime}|/2$.
-Considering the dynamical gain saturation, the corresponding value is $\kappa\approx\Lambda/4-\Lambda/3$.
+Similar to the mechanism of STDS stabilization by non-dissipative confinement along the longitudinal ($t$) coordinate Kalashnikov and Wabnitz (2022), which corresponds to an active phase mode-locking in a laser, the dissipative confinement due to synchronous pulse pumping or active amplitude mode-locking leads to a stable STDS, as well.
+Such stabilization is possible at very low values of the modulation “depth” parameter $\omega=G_{0}/\mathscr{T}^{2}\approx\Lambda/\mathscr{T}^{2}~{}10^{-6}$.
+For instance, it means that a pumping pulse, synchronized with the cavity round-trip, could be “long” compared to the STDS temporal width.For instance, the normalizations presented in the Tables 2,3 result in the dimensional $\mathscr{T}\approx 8$ ps for $\Lambda=0.1$ , 10 cm ZnS-waveguide, $|\beta_{2}|=$2000 fs2/cm, and $\omega=10^{-6}$.
+The linear stability analysis based on the variational approximation demonstrates that the stable STDS exists within confined regions of $\kappa$, $\Lambda^{\prime}$, and $\omega$ parameters.
+The numerical simulations demonstrate that the main destabilization scenario is an exponential STDS decay.
+Also, there is a tendency to STDS splitting in the time domain with the growth of $\kappa$ and $\omega$.
+For such a scenario, the “rogue wave” dynamics is characteristic.
+At variance with the case of longitudinal phase confinement Kalashnikov and Wabnitz (2022), we did not observe complicated spatial dynamics and splitting that testifies to the intensification of the mode cleaning process.
 It is important that spectral dissipation ($\tau$) is required for STDS stabilization.
 One may assume that the too-low value of this parameter cannot prevent a tendency to the STDS collapse.
-The numerical simulations demonstrated that, besides the STDS destabilization due to decay for low $|\Lambda^{\prime}|$, $\kappa\rightarrow 0\land|\Lambda^{\prime}|$, and small $\omega$ and $\tau$, there is a tendency to STDS splitting in the time domain with the growth of $\kappa$ if $\omega$ is not sufficiently high.
-That means a “transverse” confinement needs an adequate “longitudinal” one.
-Despite the case of longitudinal phase confinement [53], we did not observe complicated spatial dynamics and splitting that testifies to the intensification of the mode cleaning process.
-Although “ring-like” spatial structures and the STDS “breathing” in the transverse direction are possible in the initial stages of evolution.
 We found that the gain saturation could significantly impact the STDS dynamics.
-For the considered values of gain coefficient $G_{ss}$ there is a confined range of driving frequency $\omega$ providing the STDS stability, and this region is the broadest in the vicinity of $\kappa\approx 0.0003-0.0004$.
-The main mechanism of destabilization is STDS decay.
+For the considered values of gain coefficient $G_{ss}$ there is a confined range of driving frequency $\omega$ providing the STDS stability, and this region is the broadest in the vicinity of $\kappa\approx\Lambda/4-\Lambda/3$.
+The primary mechanism of destabilization is STDS decay.
 A possible explanation is the following.
@@
 Too large driven frequency squeezes a gain window so STDS cannot be amplified sufficiently to start a Kerr effect for spatial squeezing, and spatial confinement becomes destructive.
-Also, the dynamical gain saturation defines some maximum $\kappa$ because it decreases an effective gain, and the spatial losses lead to the STDS decay.
-STDS peak power and width oscillations appear for small $\kappa$.
+Also, the dynamic gain saturation defines some maximum $\kappa$ because it decreases an effective gain, and the spatial losses lead to the STDS decay.
+The STDS peak power and width oscillations appear with the $\kappa$-decrease.
 These oscillations can be extremely strong and accompanied by alternating between asymmetric one and two pulses.
-We must note, that the cavity round-trip period could excess essentially $2L_{w}/c$ in order to provide the STDS energy scalability [59].
-This method is practiced in high-energy passively mode-locked oscillators (e.g., see [12, 30]).
+We must note that the cavity round-trip period could excess essentially $2L_{w}/c$ to provide the STDS energy scalability Demesh et al.
+(2023).
+This method is practiced in high-energy passively mode-locked oscillators (e.g., see Südmeyer et al.
+(2008); Zhang et al.
+(2022)).
 In a waveguide oscillator, it is possible to use the Herriott cell or fiber loop for this aim.
-The last approach could be the most effective for the STDS generation in a fiber laser [53].
+The last approach could be the most effective for the STDS generation in a fiber laser Kalashnikov and Wabnitz (2022).
 In conclusion, we propose a mechanism of STDS stabilization using longitudinally and transversely graded gain in a waveguide laser with transversely graded refractive index and Kerr nonlinearity.
@@
 Numerical simulations under cylindrical symmetry conditions demonstrate the STDS’s stability within a limited region of parameters defined by net gain, pump beam width, and spectral bandwidth.
-Considering the dynamical gain saturation, we defined the main destabilization scenarios: decay and STDS oscillations asymmetrically alternating one and two pulses.
+Considering the dynamic gain saturation, we defined the main destabilization scenarios: decay and STDS oscillations asymmetrically alternating one and two pulses.
 We assume that the self-control of the transverse modes by the combination of enhanced nonlinearity, specific for the distributed nonlinear systems like waveguides composed from highly-nonlinear media, with the dissipative confinement, could provide the breakthrough in developing the new generation of high-stable and energy scalable sources of STDS.
@@
 Moreover, it could be useful for stabilizing other localized coherent structures, e.g., a weakly dissipative BEC.
-Funding
-The Norwegian Research Council projects #303347 (UNLOCK), #326503 (MIR), #32641 (Lammo-3D), and ALTA Lasers AS supported this work.
+Acknowledgments
+The authors would like to thank the Norwegian Research Council projects #303347 (UNLOCK), #326503 (MIR), #32641 (Lammo-3D), and ALTA Lasers AS supported this work.
+We are grateful to anonymous reviewers for the valuable comments that allowed us to make the essential improvements in the presentation of results.
+Appendix A Euler-Lagrange-Kantorovich equations
+The substitution of the ansatz (7) into the reduced Lagrangian (6) with a subsequent integration results in the following Euler-Lagrange equations (left-hand side of Eq(5)) for the ansatz parameters (a prime denotes $d/dZ$):
+The substitution of (7) into the right-hand side of Eq. (5) leads to the Euler-Lagrange-Kantorovich equations:
+One may obtain the equations for the evolution of the ansatz parameters from (12):
+The steady-state points of (13) give Eqs. (8,9).
+Appendix B Soliton spectrum
+To find the Fourier image of (7), we use the fact that the chirp $\chi\ll 1$ in an anomalous dispersion regime.
+This fact results from calculations of (8,9).
+Then, it is possible to use a Fourier image expansion over $\chi$:
+$\mathcal{F}\{\psi(r,t)\}(\Omega,\varpi)\approx\alpha\sqrt{\frac{2\pi\rho^{2}}{1-2i\theta\rho^{2}}}e^{-\frac{\Omega^{2}\rho^{2}}{2(1-2i\theta\rho^{2})}}\times$
+[…]
+The integral evaluation leads to:
+Appendix C Jacobian
+The calculated Jacobian for the system (13) is
+where $\theta_{0}$, $\chi_{0}$, $\rho_{0}$, $\alpha_{0}$, and $T_{0}$ are the stationary points of the system (14).